OpenStudy (lol4112):

Medal and Fan Write the ratio as a percent. In the seventh grade class, 19/25 students voted in favor of having a class picnic.

OpenStudy (anonymous):

To make this a percent, multiply the top and bottom by 4 so it is a fraction over 100. Your numerator (top number) will be the percent. \[(\frac{ 19 }{ 25 })(\frac{ 4 }{ 4 })=\frac{ 76 }{ 100 }\] So 76% of students were in favor of having a class picnic.